OpenStudy (anonymous):

If f(x) = ax^2 -12x + c and f(-3/2) = 8 is the maximum value, find the value of a and c.

OpenStudy (***[isuru]***):

ok.. it took me a while.. from the first data what is the equation u can get from the data f(-3/2) = 8

OpenStudy (anonymous):

9/4a + 18 + c = 8 9a + 4c = -40

OpenStudy (***[isuru]***):

ok... now moving on to the second .... they say that the maximum value of this graph is 8 which means that the vertex of this graph will be ( -3/2 , 8) now just forget it for a while.... if a function is written as y = ( a + b) ^2 + c then do u know how to find the minimum point of this just by looking at it ?

OpenStudy (***[isuru]***):

srry the function should be y= (ax + b) ^2 + c

OpenStudy (***[isuru]***):

if the function is y = ( ax + b)^2 + c this will take it's minimum value when the part ( ax + b)^2 becomes zero... which means that the minimum value of this function is c also if the function is a maximum y = - ( ax + b)^2 +c this will take it maximum value when the part ( ax + b)^2 becomes zero... which means here also the max value is c did u get it ?

OpenStudy (anonymous):

Yes.

OpenStudy (***[isuru]***):

so.. if u can arrange y = ax^2 -12x + c in to a format where u can write it as something like y = -( x + q)^2 +r what will be the value of r ?

OpenStudy (***[isuru]***):

 got it ?